Complete the following problem on using proper notation

kn -witrj Use the format shown m this example Use the format shown in this example: | {F)-(na) -imo mx-
Answer #1

From the given question,

a) weight of cat=1900 lb

1 lb = 0.453592 kg

1900 lb= 1900 * 0.453592 kg

=861.8248 kg

=862.0 kg

Mass=862.0 kg

weight = mg

=862.0*9.807

=8453.634

=8454 N

Weight =8454 N ( Four sig. figures)

b) 92.356- 90.35 =2.006= 2.00 ( two decimal places)

92.356 + 90.35 = 182.706 = 182.71 ( two decimal places)